please help!!
Solve the equation.
15m^2 - 9m = 0
thanks so much

Found 3 solutions by hamda, minerva, prabhjyot:
Answer by hamda(7)

15m^2 - 9m = 0
3m(5m - 3) = 0
3m = 0 , 5m - 3 = 0
m = 0 , m = 3/5 (Answer)

Answer by minerva(22)

in my opinion this is a quadratic equation.
15m^2 - 9m = 0
take common
3m(5m-3)=0
so
3m =0 or 5m-3=0
m= 0 , m=3/5

Answer by prabhjyot(165)

15m^2 - 9m = 0
m(15 m -9)=0
Equating each to zero
m=0
15m - 9=0
15m =9
m= 9/15 = 3/5
m=3/5
the solution to the equation is m=0,3/5